# Situationally-Aware Dynamics Learning

### v2

@@ V Results
We begin by introducing the evaluated scenarios in Section V-A and highlight how our approach overcomes the challenges faced by traditional approaches when used in such settings. We then describe the baselines in Section V-B. In Section V-C we cover the evaluation setup. We detail the state and action space as well as the tools for simulation and real world evaluation. In Section V-D we show that learning the situationally-aware dynamics model results in faster learning of better policies when compared to context-aware, model-based and model-free baselines. In Section V-E we present a series of ablations to better understand the contribution of the method’s components as well as the effect in the dynamics prediction with the learned hidden state representation. Finally, in Section V-F we carry out extensive experiments in simulation and the real world to conclude that our model augmented with the hidden state representation achieves better performance and safer policies in the unstructured terrain navigation task. Notably, we also summarize some emerging behavioral patterns that we observed during the real-world experiments.
Figure 3: Hardware setup used during real-world experiments.
V-A Evaluation Scenarios
@@
V-A1 Simulation:
We use the Gazebo simulator [Koenig-2004-394]. Specifically, we use Clearpath’s Inspection World [inspection_world]. It combines uneven surfaces, slopes, and varying friction properties. These conditions mimic real-world complexities, requiring the robot to adapt dynamically to changes in stability and traction while maintaining precise control, making it an ideal testbed for training and assessing our method.
Furthermore, we create four additional worlds with increasing levels of difficulty. Concretely, these are randomly generated environments created using height maps sampled using Perlin noise [perlin1985image] that provide a complex and realistic testing ground. These environments feature highly variable and irregular terrain patterns, including sudden changes in elevation, unpredictable slopes, and patches of different roughness and friction. The use of Perlin noise ensures natural-looking, non-repetitive landscapes that challenge the robot’s ability to adapt to dynamic and uneven surfaces. This makes them an effective way to assess the robustness and adaptability of our method in navigating real-world-like conditions. The generated height maps serve as input to the gazebo_terrain package for automatic Gazebo model terrain generation [gazebo_terrain]. The parameters used to generate the height maps are presented in Table I.
@@ V-A1 Simulation:
Table I: Parameters for randomly generated simulation terrains with increasing levels of difficulty. Scale adjusts slope steepness, with larger values producing smoother hills. Octaves define terrain detail, while persistence affects roughness by reducing octave amplitude. Lacunarity controls octave frequency, with higher values adding ruggedness.
V-A2 Real World:
Figure 4: Some real-world scenarios used to evaluate and compare the proposed dynamics model against the competitive baselines. (Left) Navigation across diverse terrain types, including mulch, grass, concrete, and mud, each with distinct surface properties and geometries. The final section involves a short but steep climb on mulch and grass, which presents a significant challenge due to reduced traction. (Right) Traversing a highly demanding terrain with steep ascents and descents, compounded by numerous bumps and irregularities. These features make the environment difficult to navigate safely and efficiently. Notably, our method is the only one capable of successfully completing this scenario.
Figure 4 contains two examples of the diverse real-world environments used to evaluate the situationally-aware dynamics model against the baselines, with additional scenarios detailed in Figure 14 and showcased in the accompanying video.
Given that the observation space is limited to proprioceptive odometry estimates, as described in Section V-C, critical environmental factors such as terrain type, surface irregularities, and characteristics like friction are not explicitly available to the model. Instead, these factors will need to be implicitly encoded within the hidden state representation. This representation serves to capture the unobserved dynamics that govern the robot’s interaction with its environment, allowing the model to adapt its behavior to varying conditions without direct access to external sensory data. This approach ensures that the dynamics model remains flexible and robust across diverse and unpredictable scenarios.
@@ V-B Baselines
    CaDM: Context-aware Dynamics Model, which aims to learn a “global model that can generalize across different dynamics” [lee2020context]. This is the closest baseline to our work, as they also learn a model that can detect and adapt to changes in the underlying dynamics of the agent. Additionally, although they need to learn three models (context encoder, forward and backward dynamics), they do so in parallel instead of relying on a multi-stage curriculum as other recent methods. We use an author-provided and publicly available implementation.
@@
    TD-MPC2: This is one of the leading MBRL methods at the time of writing [hansen2023td]. It is characterized for learning a latent-space dynamics model, together with an observation encoder, transition reward predictor, terminal value predictor and a return-maximizing action policy. We use an author-provided and publicly available implementation.
@@
    PETS: This method proposed by [chua2018deep] is a strong state-space dynamics learning method based on a probabilistic ensemble architecture, which in fact inspires our choice of architecture. Our implementation is based on a derivation from MBRL-Lib [Pineda2021MBRL].
@@
    CDL-CMI: This is a causally-inspired, model-based RL method that aims to increase predictive robustness by avoiding unnecessary dependencies between state variables and actions [wang2022causal]. They also learn a state abstraction for better out-of-distribution predictive performance and improved learning efficiency. We use an author-provided and publicly available implementation.
@@
    Model-Free: We train model-free policies with PPO [schulman2017proximal], SAC [haarnoja2018soft] and TD3 [fujimoto2018addressing]. This is done to present a well-rounded coverage of general RL methods and to get an idea of how recent competitive methods that also rely on auxiliary representations but with model-free policies would compare.
@@
    Dubins: This is a physics-based model [dubins-dyns] that has been used recently as an unstructured terrain navigation baseline in multiple works [siva2024self, yu2024adaptive].
V-C Experimental Setup
V-C1 State and Action Space:
@@
where $\mathbf{p}=[x,y,z]\in\mathbb{R}^{3}$ is the robot’s pose, $\mathbf{e}=[\phi,\zeta,\gamma]\in\mathbb{R}^{3}$ are the roll, pitch, and yaw angles (orientation), $\mathbf{v}=[v_{x},v_{y},v_{z}]\in\mathbb{R}^{3}$ are the linear velocity components, and $\mathbf{\Omega}=[\omega_{x},\omega_{y},\omega_{z}]\in\mathbb{R}^{3}$ are the angular velocity components. We limit the state space to the features in Eq. (27) in order to give room to our hidden state representation method to encode remaining necessary but unobserved ego- and world-characteristics.
@@
V-C2 Situation Space:
As defined in Section IV, the situation of the robot corresponds to the joint transition distribution. Therefore, the situation space $\mathbf{x}\in\mathbb{R}^{26}$ corresponds to the concatenation of state and action vectors:
@@ V-C2 Situation Space:
where the state vectors ($s_{t-1}$ and $s_{t}$) undergo preprocessing to (i) remove pose information (as the dynamics should be invariant to map-specific information) and (ii) replace orientation information (roll, pitch and yaw) with their sine and cosine transformations which enable the algorithm to uniquely identify them while avoiding discontinuities when the angles are wrapped around.
V-C3 Software:
All code was developed using ROS Noetic [Quigley09], Python 3.10.11 and PyTorch 2.3.0+cu121 [NEURIPS2019_9015]. For the sampling-based controllers we use the MBRL-Lib package [Pineda2021MBRL]. In the real-world, for localization we used FasterLIO with IMU preintegration on manifold [9718203, yu2023fast] with the Velodyne 16 LiDAR.
V-C4 Hardware:
@@
V-D Situationally-Aware Dynamics Learning
Figure 5: Training task in the Inspection World Environment [inspection_world]. The region around waypoint #3 is slippery and difficult to reach.
The situational awareness module processes the 26-dimensional situation observation (Eq. 29) as they arrive, in order to return the current representation of the hidden state $\theta\in\Theta$ symbolizing the local dynamics of the robot. This process follows the procedure outlined in Algorithm 3.
Then, the learned dynamics model $\mathcal{M}_{m}$ takes the current situation $\theta_{t}$ together with the robot’s orientation and commanded linear and angular velocities in order to predict the change in pose and orientation.
The model $\mathcal{M}_{m}$ is composed of $m=5$ probabilistic neural networks, with three dense layers with 200 neurons each and LeakyReLU non-linearities in between. The output is the mean and log-variance of a Gaussian distribution. Inputs and outputs are normalized using data statistics computed during training. At inference time, predictions are de-normalized with respect to the data statistics obtained during training. We use the Adam optimizer with default hyperparameters [kingma2014adam]. The batch size is 256 with an early-stopping threshold of 1e-3. The model $\mathcal{M}_{m}$ starts with randomly initialized parameters and is trained at the beginning of each episode with the updated replay buffer $\mathcal{R}$, analogous to the MBRL procedure shown in Algorithm 1. A summary of the hyperparameters for the situational awareness module (online hidden state representation learning) and the controller is provided in Table II.
@@ V-D Situationally-Aware Dynamics Learning
Table II: Hyperparameters used for Situational Awareness and MPPI.
V-D1 Training Results:
@@
Figure 7: Observations received by the proposed situation identification algorithm during a mission in the Inspection World. The red vertical lines mark the time when we predict a situation change happened. For ease of visualization we only present the first six features of the 26-dimensional observation, the full figure is available in Appendix A-B. Each row shows one variable’s evolution over time. The last row shows the run length probabilities indicating how likely it is that the robot stayed in the same situation (run length grows) or a situation change happened (run length becomes zero).
@@
Quantitatively evaluating the performance of online hidden state representation learning is inherently challenging for this task, as the true hidden state and its transitions are naturally unobservable. To address this, we qualitatively assess the performance of the situational awareness module through two complementary analyses. First, we examine whether the module identified meaningful situation changepoints during the robot’s operation, as illustrated in Figure 7. The results show that the model effectively segmented the local transition dynamics under varying patterns, accurately capturing and representing the changes in the dynamics. The full 26-dimensional picture is available in Appendix A-B. Second, we visualize the diversity of learned hidden state representations by mapping the 26-dimensional local transition dynamics $T_{L}$ identified during training onto the first two principal components, shown in Figure 8. This visualization reveals distinct clusters corresponding to different situations, demonstrating that each representation encodes unique patterns in the transition dynamics without collapsing into a single, generalized representation.
V-E Ablation Studies
V-E1 Learning Efficiency:
@@
We evaluate whether the advantages of our method stem from the situation information or input feature selection. We compare against other probabilistic ensembles with equivalent architectures but relying on different input features that do not include the hidden state representation. Specifically, the models include commanded linear and angular velocity as well as: (a) the robot’s heading as state input (which we call “PE-Y”); (b) all the orientation information (“PE-RPY”); (c) orientation along with linear velocities of the robot (“PE-RPY-VW”); and (d) one that also includes pose to see if additionally learning the map information results in better performance (“Full”). These models let us evaluate the contribution of our hidden state representation, as they involve multiple possible input configurations for the dynamics model but without the situation information.
@@
V-E2 Effect of the Hidden State Representations:
We use the Integrated Gradients [integrated_gradients] feature attribution method to verify that our situationally-aware dynamics model relies on the different situations to make its predictions, thus influencing the downstream planner’s control strategies. Figure 10 shows that on average, 60% of the prediction is explained by the current situation the robot is in, while the remaining input features together account for roughly 40%. This clearly highlights that the model is aware of the situation the robot is experiencing and modulates its predictions accordingly.
@@ V-E2 Effect of the Hidden State Representations:
For a more detailed examination of the situational awareness module’s behavior under controlled dynamics changes, we conduct additional experiments in a simulated environment with a manipulable UDGP. These experiments, detailed in Appendix A-A, use the Lake World Simulator to evaluate how changes in wave generation patterns affect the identification of situation transitions by an autonomous boat. The results highlight the impact of the hyperparameter $\lambda$ in segmenting the observation space and provide further insights into the relationship between UDGP variations and situation changepoints.
V-F Situationally-Aware Unstructured Terrain Navigation
@@
V-F1 Simulation Experiments:
The complementary simulation experiments were carried out in the four randomly-generated simulation environments introduced in Section V-C3, whose parameters are described in Table I. These environments feature increasing levels of difficulty with successively steeper and more rugged terrains. The experiments comprised navigation tasks from fifty randomly-generated initial and goal positions for every environment. Table III shows that our situationally-aware dynamics model had a higher success rate than the remaining methods. Nonetheless, we identify that situation switching leads the robot to re-plan under the new context which seems to slightly affect some of the path smoothness metrics (position jerk, orientation vibration and orientation jerk).
@@ V-F1 Simulation Experiments:
Control Inconsistency. In Figure 13 we measure the control inconsistency. Formally, it relates to “the error between the robot’s expected navigational behavior and its actual behavior in terms of robot poses” [siva2024self]. Our method demonstrates the best performance, followed by the model that incorporates all orientation information (PE-RPY). This result underscores the suitability of our model for sampling-based control. Specifically, our approach strikes the best balance for auto-regressive prediction tasks, such as planning with MPPI. It benefits from the lower one-step prediction error characteristic of more complex models that are harder to learn and generalize, while also offering the reduced prediction space of simpler models (e.g., Dubins), which helps mitigate compounding errors during trajectory rollout.
@@
Table III: Results on four simulation environments with increasing level of difficulty. The robot needs to traverse to fifty randomly generated initial and goal locations in each environment. We evaluate the Success Rate, Position Jerk $\frac{m}{s^{3}}]$, Orientation Vibration $\frac{rad}{s}]$ and Orientation Jerk $\frac{rad}{s^{3}}]$.
V-F2 Real World Experiments:

### v3

@@ 5 Results: Unmanned Ground Vehicle
We begin by introducing the evaluated scenarios in Section 5.1 and highlight how our approach overcomes the challenges faced by traditional approaches when used in such settings. We then describe the baselines in Section 5.2. In Section 5.3 we cover the evaluation setup. We detail the state and action space as well as the tools for simulation and real world evaluation. In Section 5.4 we show that learning the situationally-aware dynamics model results in faster learning of better policies when compared to context-aware, model-based and model-free baselines. In Section 5.5 we present a series of ablations to better understand the contribution of the method’s components as well as the effect in the dynamics prediction with the learned hidden state representation. Finally, in Section 5.6 we carry out extensive experiments in simulation and the real world to conclude that our model augmented with the hidden state representation achieves better performance and safer policies in the unstructured terrain navigation task. Notably, we also summarize some emerging behavioral patterns that we observed during the real-world experiments.
Figure 4: Hardware setup used during real world experiments.
5.1 Evaluation Scenarios
@@
5.1.1 Simulation:
We use the Gazebo simulator Koenig and Howard (2004). Specifically, we use Clearpath’s Inspection World Clearpath (2020). It combines uneven surfaces, slopes, and varying friction properties. These conditions mimic real-world complexities, requiring the robot to adapt dynamically to changes in stability and traction while maintaining precise control, making it an ideal testbed for training and assessing our method.
Furthermore, we create four additional worlds with increasing levels of difficulty. Concretely, these are randomly generated environments created using height maps sampled using Perlin noise Perlin (1985) that provide a complex and realistic testing ground. These environments feature highly variable and irregular terrain patterns, including sudden changes in elevation, unpredictable slopes, and patches of different roughness and friction. The use of Perlin noise ensures natural-looking, non-repetitive landscapes that challenge the robot’s ability to adapt to dynamic and uneven surfaces. This makes them an effective way to assess the robustness and adaptability of our method in navigating real-world-like conditions. The generated height maps serve as input to the gazebo_terrain package for automatic Gazebo model terrain generation Verbryke (2021). The parameters used to generate the height maps are presented in Table 1.
@@ 5.1.1 Simulation:
Table 1: Parameters for randomly generated simulation terrains with increasing levels of difficulty. Scale adjusts slope steepness, with larger values producing smoother hills. Octaves define terrain detail, while persistence affects roughness by reducing octave amplitude. Lacunarity controls octave frequency, with higher values adding ruggedness.
5.1.2 Real World:
Figure 5: Some real-world scenarios used to evaluate and compare the proposed dynamics model against the competitive baselines. (Left) Navigation across diverse terrain types, including mulch, grass, concrete, and mud, each with distinct surface properties and geometries. The final section involves a short but steep climb on mulch and grass, which presents a significant challenge due to reduced traction. (Right) Traversing a highly demanding terrain with steep ascents and descents, compounded by numerous hidden bumps and irregularities. These features make the environment difficult to navigate safely and efficiently. Notably, our method is the only one capable of successfully completing this scenario.
Figure 5 contains two examples of the diverse real-world environments used to evaluate the situationally-aware dynamics model against the baselines, with additional scenarios detailed in Figure 15 and showcased in the accompanying videos.
Given that the observation space is limited to proprioceptive odometry estimates, as described in Section 5.3, critical environmental factors such as terrain type, surface irregularities, and characteristics like friction are not explicitly available to the model. Instead, these factors will need to be implicitly encoded within the hidden state representation. This representation serves to capture the unobserved dynamics that govern the robot’s interaction with its environment, allowing the model to adapt its behavior to varying conditions without direct access to external sensory data.
@@ 5.2 Baselines
    CaDM: Context-aware Dynamics Model, which aims to learn a “global model that can generalize across different dynamics” Lee et al. (2020b). This is the closest baseline to our work, as they also learn a model that can detect and adapt to changes in the underlying dynamics of the agent. Additionally, although they need to learn three models (context encoder, forward and backward dynamics), they do so in parallel instead of relying on a multi-stage curriculum as other recent methods. We use an author-provided and publicly available implementation.
@@
    TD-MPC2: This is one of the leading MBRL methods at the time of writing Hansen et al. (2024). It is characterized for learning a latent-space dynamics model, together with an observation encoder, transition reward predictor, terminal value predictor and a return-maximizing action policy. We use an author-provided and publicly available implementation.
@@
    PETS: This method proposed by Chua et al. (2018) is a strong state-space dynamics learning method based on a probabilistic ensemble architecture, which in fact inspires our choice of architecture. Our implementation is based on a derivation from MBRL-Lib Pineda et al. (2021).
@@
    CDL-CMI: This is a causally-inspired, model-based RL method that aims to increase predictive robustness by avoiding unnecessary dependencies between state variables and actions Wang et al. (2022). They also learn a state abstraction for better out-of-distribution predictive performance and improved learning efficiency. We use an author-provided and publicly available implementation.
@@
    Model-Free: We train model-free policies with PPO Schulman et al. (2017), SAC Haarnoja et al. (2018) and TD3 Fujimoto et al. (2018). This is done to present a well-rounded coverage of general RL methods and to get an idea of how recent competitive methods that also rely on auxiliary representations but with model-free policies would compare.
@@
    Dubins: This is a physics-based model Dubins (1957) that has been used recently as an unstructured terrain navigation baseline in multiple works Siva et al. (2024); Yu et al. (2024).
5.3 Experimental Setup
5.3.1 State and Action Space:
@@
where $\mathbf{p}=[x,y,z]\in\mathbb{R}^{3}$ is the robot’s position, $\mathbf{e}=[\phi,\zeta,\gamma]\in\mathbb{R}^{3}$ are the roll, pitch, and yaw angles (orientation), $\mathbf{v}=[v_{x},v_{y},v_{z}]\in\mathbb{R}^{3}$ are the linear velocity components, and $\mathbf{\Omega}=[\omega_{x},\omega_{y},\omega_{z}]\in\mathbb{R}^{3}$ are the angular velocity components. We limit the state space to the elements in Eq. (27) in order to give room to our hidden state representation method to encode remaining necessary but unobserved ego- and world-characteristics.
@@
5.3.2 Situation Space:
As defined in Section 4, the situation of the robot corresponds to the joint transition distribution. Therefore, the situation space $\mathbf{x}\in\mathbb{R}^{26}$ corresponds to the concatenation of state and action vectors:
@@ 5.3.2 Situation Space:
where the state vectors ($s_{t-1}$ and $s_{t}$) undergo preprocessing to (i) remove position information (as the dynamics should be invariant to map-specific information) and (ii) replace orientation information (roll, pitch and yaw) with their sine and cosine transformations which enable the algorithm to uniquely identify them while avoiding discontinuities when the angles are wrapped around.
5.3.3 Software:
All code was developed using ROS Noetic Quigley et al. (2009), Python 3.10.11 and PyTorch 2.3.0+cu121 Paszke et al. (2019). For the sampling-based controllers we use the MBRL-Lib package Pineda et al. (2021). In the real-world, for localization we used FasterLIO with IMU preintegration on manifold Bai et al. (2022); Yu et al. (2023) with the Velodyne 16 LiDAR.
5.3.4 Hardware:
@@
5.4 Situationally-Aware Dynamics Learning
Figure 6: Training task in the Inspection World Environment Clearpath (2020). The region around waypoint #3 is slippery and difficult to reach.
The situational awareness module processes the 26-dimensional situation observation (Eq. 29) as they arrive, in order to return the current representation of the hidden state $\theta\in\Theta$ symbolizing the local dynamics of the robot. This process follows the procedure outlined in Algorithm 3. Then, the learned dynamics model $\mathcal{M}_{m}$ takes the current situation $\theta_{t}$ together with the robot’s orientation and commanded linear and angular velocities in order to predict the change in position and orientation.
The model $\mathcal{M}_{m}$ is composed of $m=5$ probabilistic neural networks, with three dense layers with 200 neurons each and LeakyReLU non-linearities in between. The output is the mean and log-variance of a Gaussian distribution. Inputs and outputs are normalized using data statistics computed during training. At inference time, predictions are de-normalized with respect to the data statistics obtained during training. We use the Adam optimizer with default hyperparameters Kingma and Ba (2014). The batch size is 256 with an early-stopping threshold of 1e-3. The model $\mathcal{M}_{m}$ starts with randomly initialized parameters and is trained at the beginning of each episode with the updated replay buffer $\mathcal{R}$, analogous to the MBRL procedure shown in Algorithm 1. That is, we follow the active training procedure described in Section 4.1.1. A summary of the hyperparameters for the situational awareness module (online hidden state representation learning) and the controller is provided in Table 2.
@@ 5.4 Situationally-Aware Dynamics Learning
Table 2: Hyperparameters used for Situational Awareness and MPPI in the UGV unstructured terrain navigation task.
5.4.1 Training Results:
@@
Figure 8: Observations received by the proposed situation identification algorithm during a mission in the Inspection World. The red vertical lines mark the time when we predict a situation change happened. For ease of visualization we only present the first six features of the 26-dimensional observation, the full figure is available in Appendix B.2. Each row shows one variable’s evolution over time. The last row shows the run length probabilities indicating how likely it is that the robot stayed in the same situation (run length grows) or a situation change happened (run length becomes zero).
@@
Quantitatively evaluating the performance of online hidden state representation learning is inherently challenging for this task, as the true hidden state and its transitions are naturally unobservable. To address this, we qualitatively assess the performance of the situational awareness module through two complementary analyses. First, we examine whether the module identified meaningful situation changepoints during the robot’s operation, as illustrated in Figure 8. The results show that the model effectively segmented the local transition dynamics under varying patterns, accurately capturing and representing the changes in the dynamics. The full 26-dimensional picture is available in Appendix B.2. Second, we visualize the diversity of learned hidden state representations by mapping the 26-dimensional local transition dynamics $T_{L}$ identified during training onto the first two principal components, shown in Figure 9. This visualization reveals distinct clusters corresponding to different situations, demonstrating that each representation encodes unique patterns in the transition dynamics without collapsing into a single, generalized representation.
5.5 Ablation Studies
5.5.1 Learning Efficiency:
@@
We evaluate whether the advantages of our method stem from the situation information or input feature selection. We compare against other probabilistic ensembles with equivalent architectures but relying on different input features that do not include the hidden state representation. Specifically, the models include commanded linear and angular velocity as well as: (a) the robot’s heading as state input (which we call “PE-Y”); (b) all the orientation information (“PE-RPY”); (c) orientation along with linear velocities of the robot (“PE-RPY-VW”); and (d) one that also includes position to see if additionally learning the map information results in better performance (“Full”). These models let us evaluate the contribution of our hidden state representation, as they involve multiple possible input configurations for the dynamics model but without the situation information.
@@
5.5.2 Effect of the Hidden State Representations:
We use the Integrated Gradients Sundararajan et al. (2017) feature attribution method to verify that our situationally-aware dynamics model relies on the different situations to make its predictions, thus influencing the downstream planner’s control strategies. Figure 11 shows that on average, 60% of the prediction is explained by the current situation the robot is in, while the remaining input features together account for roughly 40%. This clearly highlights that the model is aware of the situation the robot is experiencing and modulates its predictions accordingly.
@@ 5.5.2 Effect of the Hidden State Representations:
For a more detailed examination of the situational awareness module’s behavior under controlled dynamics changes, we conduct additional experiments in a simulated environment with a manipulable UDGP. These experiments, detailed in Appendix B.1, use the Lake World Simulator to evaluate how changes in wave generation patterns affect the identification of situation transitions by an autonomous boat. The results highlight the impact of the hyperparameter $\lambda$ in segmenting the observation space and provide further insights into the relationship between UDGP variations and situation changepoints.
5.6 Situationally-Aware Unstructured Terrain Navigation
@@
5.6.1 Simulation Experiments:
The complementary simulation experiments were carried out in the four randomly-generated simulation environments introduced in Section 5.1.1, whose parameters are described in Table 1. These environments feature increasing levels of difficulty with successively steeper and more rugged terrains. The experiments comprised navigation tasks from fifty randomly-generated initial and goal positions for every environment. Table 3 shows that our situationally-aware dynamics model had a higher success rate than the remaining methods. Nonetheless, we identify that situation switching leads the robot to re-plan under the new context which seems to slightly affect some of the path smoothness metrics (position jerk, orientation vibration and orientation jerk).
@@ 5.6.1 Simulation Experiments:
Control Inconsistency. In Figure 14 we measure the control inconsistency. Formally, it relates to “the error between the robot’s expected navigational behavior and its actual behavior in terms of robot poses” Siva et al. (2024). Our method demonstrates the best performance, followed by the model that incorporates all orientation information (PE-RPY).
This result underscores the suitability of our model for sampling-based control. Specifically, our approach strikes the best balance for auto-regressive prediction tasks, such as planning with MPPI. It benefits from the lower one-step prediction error characteristic of more complex models that are harder to learn and generalize, while also offering the reduced prediction space of simpler models (e.g., Dubins), which helps mitigate compounding errors during trajectory rollout.
@@
Table 3: Results on four simulation environments with increasing level of difficulty. The robot needs to traverse to fifty randomly generated initial and goal locations in each environment. We evaluate the Success Rate, Position Jerk $\frac{m}{s^{3}}]$, Orientation Vibration $\frac{rad}{s}]$ and Orientation Jerk $\frac{rad}{s^{3}}]$.
5.6.2 Real World Experiments: